OpenStudy (anonymous):

(3x + 2) (2x - 5) = ax² + kx + n . What is the value of a - n + k ?

OpenStudy (anonymous):

do you know what they are asking or not?

OpenStudy (anonymous):

This is a quadratic equation. To find n we multiply the last terms in each bracket together (2 times -5) = -10 To find ‘a’ we multiply the first terms in each bracket together (3x times 2x) = 6x2, So a = 6. Thus a - n = 16 To find k we multiply the first term in the first bracket by the second term in the second bracket and add the result to the product of the second term in the first bracket and the first term in the second. (-15x + 4x) = -11x. So k = -11 16 - 11 = 5

OpenStudy (anonymous):

@ankita808 do u agree with my sollution